# Ultrafast broadband soliton microcomb laser

### v1

@@ Introduction
Figure 4: Optical linewidth of the soliton microcomb laser. (a) Frequency noise spectrum of individual comb lines and the overall comb for the 1.6-THz soliton state. The overall comb has frequency noise floor lower than individual comb lines likely because of detector noises that limit the characterizaton of the latter due to their lower powers. (b) Spectral distribution of intrinsic optical linewidth for individual comb lines $\Delta\nu_{m}$, along with the soliton comb spectrum. (c) Comparison of intrinsic optical linewidth and optical efficiency across different on-chip soliton microcomb sources. Complete performance metrics and definitions are provided in Table S3 of SI.
Frequency noise of an OFC determines its coherence length which underlies crucially many applications in communications, sensing, spectroscopy, and metrology. To quantify the intrinsic linewidth of our soliton microcomb laser, we employed the correlated self‑heterodyne technique 46 to measure the Lorentzian linewidth of individual comb teeth as well as the overall comb for the 1.6-THz soliton comb state at 120 mA (Fig. S3).
Fig. 4 shows the recorded frequency noise spectra and the corresponding comb linewidths. The intrinsic linewidth distribution across the spectrum exhibits a distinctive quadratic profile as shown in Fig. 4a. This parabolic dependence of individual comb‑line linewidths further corroborates the nature of mode-locked solitons, in which the frequency of a comb tooth is given by $\nu_{m}=\nu_{\mathrm{c}}+m\,\nu_{\mathrm{rep}}$ where $\nu_{\mathrm{c}}$ and $\nu_{\mathrm{rep}}$ are the center comb‑line frequency and the repetition frequency, respectively, and $m$ is the mode index relative to the center comb mode, leading to a $m^{2}$ dependence of individual comb mode linewidth 36, 47. Fig. 4a shows that the overall soliton comb exhibits a white frequency noise floor of 17 $\rm Hz^{2}/Hz$}, which corresponds to an intrinsic Lorentzian linewidth as low as only 53 Hz, a value that is even comparable with state-of-the-art on-chip single-frequency CW lasers 48. The superior linewidth performance of our soliton laser results directly from the nature of ultralow-loss TFLN external laser cavity structure that combines elegantly with an embedded high-Q microresonator for efficient mode locking. Fig. 4b shows that the spectral distribution of comb tooth linewidths displays an asymmetry about the central mode. The 1598 nm tooth demonstrates a significantly narrower Lorentzian linewidth compared to the 1540 nm mode, despite both being equidistant ($m=\pm 2$) from the central mode. We attribute this asymmetry to stimulated Raman scattering (SRS) within the soliton, a phenomenon previously observed in DKS microcomb systems 21.
@@ Discussion
In addition to comb linewidth, another crucial metric underlying microcomb applications is power efficiency. A proper metric to compare across different on-chip comb sources is optical efficiency defined as the efficiency of converting optical power to comb 49. Fig. 4c shows the performance of representative state-of-the-art on-chip comb sources. Apparently, SMLLs exhibit near-unity optical efficiency inherent to their laser nature, yet suffer from poor linewidth due to high intracavity losses. DKS’s linewidth relies on the pump laser and could achieve fairly narrow linewidth with an ultra-coherent pump laser, while its optical efficiency remains relatively low compared with SMLLs. Fig. 4c shows that our laser device combines nicely the advantages of these two systems. On one hand, it exhibits near-unity optical efficiency same as SMLLs thanks to its essential mode-locked laser nature. At the same time, it has the smallest linewidth ever demonstrated for on-chip soliton microcombs including DKSs and SMLLs 7, 8, 36, even surpassing DKS that is pumped with an ultra-narrow-linewidth fiber-based external cavity laser 50.
To conclude, we have demonstrated the first fully integrated, background-free, electrically pumped soliton microcomb laser via InP/TFLN hybrid integration. The chip-scale mode-locked laser achieves remarkable performance with a broad 3-dB bandwidth of >>3.4 THz, an ultrahigh repetition frequency up to 3 THz, and an ultra-narrow linewidth down to 53 Hz. Our approach dramatically simplifies bright-soliton generation with stable turn-key operation, near-unity optical efficiency, superior long-term stability, and remarkably low soliton generation threshold (1 V, 75 mA) that could be powered with a standard alkaline battery. Therefore, the demonstrated soliton laser represents a paradigm shift in chip-scale frequency comb technology, providing a practical path for transitioning soliton microcomb technology from controlled laboratory environments to diverse real-world applications. Moreover, together with the strong Pockels effect and quadratic nonlinearity of TFLN, our laser now opens up a great avenue towards full on-chip integration of soliton comb laser with multiple electro-optic and frequency conversion functionalities that would open the door for a wide range of applications such as frequency self-referencing, frequency synthesis/division, high-speed photonic signal processing, comb waveband conversion, among many others, all directly integrated on a single chip.
@@ Acknowledgements
We thank Dr. Daniel Kane of MesaPhotonics for helpful discussions about the FROG traces. We also thank Dr. Wuxiucheng Wang and Yongchao Liu for their help with setup and current measurement. This work is supported in part by the Defense Advanced Research Projects Agency (DARPA) NaPSAC program (Agreement No. N660012424007) and the National Science Foundation (NSF) (OMA-2138174, ECCS-2231036, OSI-2329017). This work was performed in part at the Cornell NanoScale Facility, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation (Grant NNCI-2025233).

### v2

@@ Introduction
Figure 4: Optical linewidth of the soliton microcomb laser. (a) Frequency noise spectrum of individual comb lines and the overall comb for the 1.6-THz soliton state. The overall comb has frequency noise floor lower than individual comb lines likely because of detector noises that limit the characterizaton of the latter due to their lower powers. (b) Spectral distribution of intrinsic optical linewidth for individual comb lines $\Delta\nu_{m}$, along with the soliton comb spectrum. (c) Comparison of intrinsic optical linewidth and optical efficiency across different on-chip ultrafast comb sources. Complete performance metrics and definitions are provided in Table S2 of SI.
Another significant challenge hindering the deployment of conventional on-chip MLLs is their limited coherence, a property that is critical for applications in communications, sensing, spectroscopy, and metrology. Most existing on-chip MLLs exhibit intrinsic optical linewidths in the sub-MHz range range, a limitation largely attributed to the high optical loss inherent to III-V materials17, 18. To quantify the intrinsic linewidth of the soliton laser, we employed the correlated self‑heterodyne technique to measure the Lorentzian linewidth of individual comb teeth as well as the overall comb for the 1.6-THz soliton comb state at 120 mA (Fig. S3).
Fig. 4 shows the recorded frequency noise spectra and corresponding comb linewidths. As shown in Fig. 4a, the overall soliton comb exhibits a white frequency noise floor of 17 $\rm Hz^{2}/Hz$}, corresponding to an intrinsic Lorentzian linewidth as low as 53 Hz. This value represents an improvement of at least three orders of magnitude over the best current chip-scale MLLs (Fig. 4c) and is comparable to advanced on-chip free-running CW lasers44. The superior linewidth performance of the soliton laser results directly from the nature of ultralow-loss TFLN external laser cavity structure that combines elegantly with an embedded high-Q microresonator for efficient mode locking.
The intrinsic linewidth distribution across the spectrum exhibits a distinctive quadratic profile (Fig. 4b). This parabolic dependence of individual comb-line linewidths further corroborates the mode-locked soliton nature, specifically the ‘elastic-tape’ behavior characteristic of passively locked solitons45, 46. Because comb modes are grown via FWM symmetrically around the center, consequently, comb lines further away from the center are more strongly affected by repetition rate jitter, leading to a squared dependence of linewidth on the mode number. Fig. 4b shows that the spectral distribution of comb tooth linewidths displays an asymmetry about the central mode. The 1598 nm tooth demonstrates a significantly narrower Lorentzian linewidth compared to the 1540 nm mode, despite both being equidistant ($m=\pm 2$) from the central mode. We attribute this asymmetry to stimulated Raman scattering (SRS) within the soliton, a phenomenon previously observed in DKS microcomb systems 47.
@@ Discussion
To conclude, we have demonstrated the first on-chip femtosecond laser, enabled by photorefraction-assisted soliton mode-locking. This device achieves pulse durations shorter than 90 fs with repetition rates up to 3 THz, while simultaneously exhibiting an intrinsic linewidth of 53 Hz. Our approach dramatically simplifies on-chip bright-soliton generation, offering stable turn-key operation, near-unity optical efficiency, superior long-term stability, and a remarkably low generation threshold (1 V, 75 mA) compatible with a standard alkaline battery. This soliton laser elegantly synthesizes the superior pulse duration and coherence of DKS microcombs with the operational simplicity and high efficiency of conventional on-chip MLLs (Fig. 4c). The demonstrated soliton laser represents a fundamental advance in chip-scale frequency comb and pulse source technology, providing a compact and efficient solution that unlocks femtosecond pulse systems from controlled laboratory environments for diverse, real-world applications.
In the future, further integration with electrical tuning could possibly help fundamental mode-locking and high-speed soliton modulation48. By leveraging the rich optical properties of TFLN, particularly its quadratic nonlinearity and Pockels effect49, 50, this work opens a significant avenue toward the full on-chip integration of soliton comb lasers with multiple electro-optic and frequency conversion functionalities. This could accelerate various applications, such as self-referencing, frequency synthesis, high-speed photonic signal processing, and comb waveband conversion, all to be directly integrated on a single chip.
@@ Acknowledgements
We thank Dr. Daniel Kane of MesaPhotonics for helpful discussions about the FROG traces. We also thank Dr. Wuxiucheng Wang and Yongchao Liu for their help with setup and current measurement. We thank Dr. Xiyuan Lu from NIST for the discussion about this work. This work is supported in part by the Defense Advanced Research Projects Agency (DARPA) NaPSAC program (Agreement No. N660012424007) and the National Science Foundation (NSF) (OMA-2138174, ECCS-2231036, OSI-2329017). This work was performed in part at the Cornell NanoScale Facility, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation (Grant NNCI-2025233).